The ratio of the numerator to the denominator of a certain fraction is three to five. If two is added to the numerator and five is subtracted from the denominator, the new fraction reduces to four-fifths. Find the original fraction.
Answer by Alan3354(69443)

The ratio of the numerator to the denominator of a certain fraction is three to five. If two is added to the numerator and five is subtracted from the denominator, the new fraction reduces to four-fifths. Find the original fraction.
----------------
Original is 3x/5x
---
(3x+2)/(5x-5) = 4/5
5(3x+2) = 4(5x-5)
15x+10 = 20x-20
30 = 5x
x = 6
--> 18/30